John is twice Joan's age. 5 years ago he was 7 times her age. How old will each be next year?
Answer by TimothyLamb(4379)

x = john age now
y = joan age now
---
x = 2y
x - 5 = 7(y - 5)
---
x = 2y
x - 5 = 7y - 35
---
x - 2y = 0
x - 7y = -30
---
put the system of linear equations into standard form
---
x - 2y = 0
x - 7y = -30
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 12
y = 6
---
answer:
x = john age next year = 13
y = joan age next year 7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php